A large positively-charged object with charge q = 3.25 ?C is brought near a negatively-charged plastic ball suspended from a string of negligible mass. The suspended ball has a charge of q
Answer #1

force on the suspended ball, F = k q1 q2 / r2

= 9 * 109 * 3.25 * 10-6 * 54.3 * 10-9 / 0.2152

= 0.0344 N

let t be the rension in the string.

then T sin theta = 0.0344 ---------------------(1)

T cos theta = 16.5 * 10-3 * 9.81 = 0.162 N ----------------------(2)

dividing eqn(1) by eqn (2)

tan theta = 0.212

theta = 11.97 degree